OpenStudy (poetryprincess):

Help Solving Exponents? ~Confused To The Max~ 1. m^2n^0 2. -3f^3g^-1

OpenStudy (anonymous):

Cool expressions

OpenStudy (anonymous):

\(n^0=1\) so the first one is \(m^2\)

OpenStudy (anonymous):

\[-3f^3g^{-1}=-\frac{3f^3}{g}\] not sure what else they want you to do

OpenStudy (anonymous):

@satellite73 I think the first one is m^2n^0. If you take 2n^0 this is equivalent to 1, so you would get m^1.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I was writing how to do the second one and I just realized sat already did >_>

OpenStudy (anonymous):

\[x ^{-n}=\frac{ 1 }{ x^n }\]

OpenStudy (poetryprincess):

Thanks for the help ya'll.